# The flexible Gumbel distribution: A new model for inference about the mode

### v2

@@ 4 Simulation study
In the first experiment, referred to as (E1) in the sequel, we drew a random sample of size $n\in\{100,200\}$ from an FG distribution with $\theta=0$, $\sigma_{1}=1$, $\sigma_{2}=5$, and $w=0.5$. Based on each simulated data set, we estimated $\mathbf{\Omega}$ by applying the ECM algorithm and the Metropolis-within-Gibbs algorithm. The former algorithm produced the MLE of $\mathbf{\Omega}$, and we used the median of the posterior distribution of $\mathbf{\Omega}$ at convergence of the latter algorithm as another point estimate of $\mathbf{\Omega}$. Table 1 presents summary statistics of these estimates of $\mathbf{\Omega}$ and estimates of the corresponding standard deviation across 1000 Monte Carlo replicates.
Table 1: Frequentist and Bayesian inference results in experiment (E1) across 1000 Monte Carlo replicates. Here, point.est stands for the average of 1000 point estimates for each parameter from each method, $\widehat{\text{s.d.}}$ stands for the average of the corresponding 1000 estimated standard deviations, and s.d. refers to the empirical standard deviation of the 1000 point estimates from each method. Numbers in parentheses are 100×100$\times$100 × Monte Carlo standard errors associated with the averages.
@@
According to Table 1, all estimates for parameters in $\mathbf{\Omega}$ are reasonably close to the truth. A closer inspection on the reported empirical mean of these estimates along with their empirical standard error suggests that, when $n=100$, the Bayesian method may slightly underestimate $\sigma_{2}$, the larger of the two scale parameters of FG. We believe that this is due to the inverse gamma prior imposed on the scale parameters that is sharply peaked near zero, and thus the posterior median of the larger scale parameter tends to be pulled downwards when the sample size is not large. As the sample size increases to $200$, this trend of underestimation appears to diminish. The empirical means of the standard deviation estimates from both methods are close to the corresponding empirical standard deviations, which indicate that the variability of a point estimator is accurately estimated, whether it is based on the sandwich variance estimator in the frequentist framework, or based on the posterior sampling in the Bayesian framework. In summary, the methods proposed in Section 3 under both frameworks provide reliable inference for $\mathbf{\Omega}$ along with accurate uncertainty assessment of the point estimators when data arise from an FG distribution.
@@
Among all existing mixture distributions, normal mixtures probably have the longest history and are most referenced in the literature. In another experiment, we compared the model fitting of normal mixture with that of FG when data arise from three heavy-tailed distributions: (E2) Laplace with the location parameter equal to zero and the scale parameter equal to 2; (E3) a mixture of two Gumbel distributions for the maximum, with a common mode at zero, scale parameters in the two components equal to 2 and 6, respectively, and the mixing proportion equal to 0.5; (E4) a $t$ distribution with degrees of freedom equal to 5. From each of the three distributions in (E2)–(E4), we generated a random sample of size $n=200$, following which we fit a two-component normal mixture model via the EM algorithm implemented using the R package mixtools, and also fit an FG model via the two algorithms described in Section 3. This model fitting exercise was repeated for 1000 Monte Carlo replicates under each of (E2)–(E4).
@@
We used an empirical version of the Kullback-Leibler divergence as the metric to assess the quality of modeling fitting. We denote the true density function as $p(\cdot)$, and let $\hat{p}(\cdot)$ be a generic estimated density resulting from one of the three considered model fitting strategies. Under each setting in (E2)–(E4), a random sample of size 50000, $(x_{1},\ldots,x_{50000})$, were generated from the true distribution, and an empirical version of the Kullback-Leibler divergence from $\hat{p}(\cdot)$ to $p(\cdot)$ is given by $D_{\hbox{\mbox{KL}}}=(1/50000)\sum_{i=1}^{50000}\log(p(x_{i})/\hat{p}(x_{i}))$. Figure 1 shows the boxplots of $D_{\hbox{\mbox{KL}}}$ across 1000 Monte Carlo replicates corresponding to each model fitting scheme under (E2)–(E4).
@@
Judging from Figure 1, the FG distribution clearly outperform the normal mixture when fitting data from any of the three heavy-tailed distributions in (E2)–(E4), and results from the frequentist method are comparable with those from the Bayesian method for fitting an FG model. When implementing the ECM algorithm for fitting the FG model and the EM algorithm for fitting the normal mixture, we set a maximum number of iterations at 1000. Our ECM algorithm always converged in the simulation, i.e., converged to a stationary point within 1000 iterations. But the EM algorithm for fitting a normal mixture often had trouble achieving that, with more difficulty when data come from a heavier-tailed distribution. More specifically, under (E4), which has the highest kurtosis (equal to 9) among the three settings, the EM algorithm failed to converge in 59.9% of all Monte Carlo replicates; under (E2), which has the second highest kurtosis (equal to 6), it failed to converge in 6.7% of the replicates. Results associated with the normal mixture from these failing replicates were not included when producing the boxplots in Figure 1. In conclusion, the FG distribution is more suitable for symmetric or asymmetric heavy-tailed data than the normal mixture distribution.
@@ 5 An application in hydrology
Daily maximum water elevation changes of a waterbody, such as ocean, lake, and wetland, are of interest in hydrologic research. These changes may be close to zero in most days, but can be extremely large or small under extreme weather. From National Water Information System (https://waterdata.usgs.gov/), we downloaded water elevation data for Lake Murray near Columbia, South Carolina, United States, recorded from September 18, 2020 to September 18, 2021. The water elevation change of a given day was calculated by contrasting the maximum elevation and the minimum elevation on that day, returning a positive (negative) value if the maximum record of the day comes after (before) the minimum record on the same day. We fit the FG distribution to the resultant data with $n=366$ records using the frequentist method and the Bayesian method, with results presented in Table 2. The two inference methods produced very similar estimates for most parameters, although small differences were observed. For example, one would estimate the mode of daily maximum water elevation change to be $-0.795$ feet based on the frequentist method, but estimate it to be $-0.486$ feet using the Bayesian method. The discrepancy between these two mode estimates is minimal considering that the daily maximum water elevation changes range from $-38$ feet to 49.4 feet within this one-year period. In fact, taking into account the uncertainty in these point estimates, we do not interpret any of these differences as statistically significant because a parameter estimate from one method always falls in the interval estimate for the same parameter from the other method according to Table 2. Using parameter estimates in Table 2 in the aforementioned R Shiny app, we obtained an estimated skewness of $-0.102$ and an estimated kurtosis of 6.384 based on the frequentist inference results, whereas the Bayes inference yielded an estimated skeweness of 0.058 and an estimated kurtosis of 6.074. Combining these two sets of results, we concluded that the underlying distribution of daily maximum water elevation change may be nearly symmetry, with outliers on both tails that cause tails heavier than that of a Gumbel distribution.
@@
$\theta$0.795 0.796 -2.355 0.764 -0.486 0.694 -1.679 0.973
$\sigma_{1}$ 5.186 0.541 4.124 6.247 5.399 0.651 4.534 6.916
$\sigma_{2}$ 6.237 1.735 2.836 9.638 5.734 1.031 4.390 8.029
$w$ 0.698 0.169 0.367 1.029 0.630 0.141 0.329 0.847
Figure 2 presents the estimated density functions from these two methods, in contrast with the estimated density curve resulting from fitting the data to a two-component normal mixture, and a kernel density estimate using a Gaussian kernel with the bandwidth selected according to the method proposed by Sheather and Jones [1991]. The last estimate is fully nonparametric and served as a benchmark against which the other three density estimates were assessed graphically. The kernel density estimate is more flexible at describing varying tail behaviors, but such flexibility comes at the cost of statistical efficiency and interpretability. With the wiggly tails evident in Figure 2 for this estimate, we suspected certain level of overfitting of the kernel density estimate. This often happens to kernel-based estimation of a function around a region where data are scarce, with a bandwidth not large enough for the region. Between the two FG density estimates, the difference is almost negligible. They both track the kernel density estimate closely over a wide range of the support around the mode. The mode of the estimated normal mixture density is close to the other three mode estimates, but the tails are much lighter than those of the other three estimated densities.
@@
Besides comparing the three parametric density estimates pictorially, we also used the Monte-Carlo based one-sample Kolmogorov–Smirnov test to assess the goodness of fit. The $p$-values from this test are 0.223, 0.312, and 0.106 for the frequentist FG density estimate, the Bayesian FG density estimate, and the estimated normal mixture density, respectively. Although none of the $p$-values are low enough to indicate lack of fit (at significance level 0.05 for example), the $p$-value associated with the normal mixture is much lower than those for FG. This provides quantitative evidence that an FG distribution fits the current data better than a normal mixture. It is also worth noting that the Kolmogorov–Smirnov test is known to have low power to detect deviations from a posited distribution that occur in the tails [Mason and Schuenemeyer, 1983]. This may explain the above-0.05 $p$-value for the normal mixture fit of the data even though the tail of this posited distribution may be too thin for the current data.
We used Stan to implement the Bayesian inference for the Lake Murray data, and the code and posterior output are given in the Appendix. The output provided there indicates that our MCMC chain has converged (see the Rhat statistics). The JAGS code for fitting the FG distribution is also given in the Appendix.
@@ 6 An application in criminology
Table 4: Mean regression model based on the normal distribution fitted to the crime data. Besides parameter estimates (under point.est) and the estimated standard deviations of these parameter estimates (under $\widehat{\mbox{s.d.}}$), 95% confidence intervals of the parameters
@@
With the location parameter $\theta$ signified in the FG distribution as the mode, it is straightforward to formulate a modal regression model that explores the relationship between the response variable and predictors. To demonstrate the formulation of a modal regression model based on the FG distribution, we analyze a data set from Agresti et al. [2021] in the area of criminology. This data set contains the percentage of college education, poverty percentage, metropolitan rate, and murder rate for the 50 states in the United States and the District of Columbia from year 2003. The poverty percentage is the percentage of the residents with income below the poverty level; the metropolitan rate is defined as the percentage of population living in the metropolitan area; and the murder rate is the annual number of murders per $100,000$ people in the population.
@@
Table 3 shows the inference results from the modal regression model, and Table 4 presents the inference results from the mean regression model. At $5\%$ significance level, both frequentist and Bayesian modal regression analyses confirm that there exists a negative association between the percentage of college education and the murder rate, as well as a positive association between the metropolitan rate and the murder rate. In contrast, according to the inferred mean regression model, there is a positive association between the percentage of college education and the murder rate. Such claimed positive association is intuitively difficult to justify and contradicts with many published results in criminology [Hjalmarsson and Lochner, 2012, Lochner, 2020].
The scatter plot of the data in Figure 3 can shed some light on why one reaches to such a drastically different conclusion on a covariate effect when mean regression is considered in place of modal regression. As shown in Figure 3, the exists an obvious outlier, District of Columbia (DC), in panels of the first row of the scatter plot matrix for instance. Mean regression reacts to this one extreme outlier by inflating the covariate effect associated with the percentage of college education in the inferred mean regression function. Thanks to the heavy-tailed feature of the FG distribution, modal regression based on this distribution is robust to outliers, which strives to capture data features suggested by majority of the data and is not distracted by the extreme outlier when inferring covariate effects in this application.
Figure 3: Scatter plot matrix of the crime data.

### v3

@@ 4 Simulation study
In the first experiment, denoted as (E1) hereafter, we considered two FG distributions as true data generating mechanisms, $\text{FG}(\theta=1,\sigma_{1}=1,\sigma_{2}=1,w=0.4)$ and $\text{FG}(\theta=0,\sigma_{1}=1,\sigma_{2}=5,w=0.5)$. This design creates two FG distributions with the second one more skewed and variable than the first. Based on a random sample of size $n=50$ from the first FG distribution, we estimated $\mathbf{\Omega}$ by applying the ECM algorithm and the Metropolis-within-Gibbs algorithm. Similarly, based on a random sample of size $n\in\{100,200\}$, we implemented the two algorithms to estimate $\mathbf{\Omega}$. The former algorithm produced the MLE of $\mathbf{\Omega}$, and we used the median of the posterior distribution of $\mathbf{\Omega}$ at convergence of the latter algorithm as another point estimate of $\mathbf{\Omega}$. Table 1 presents summary statistics of these estimates of $\mathbf{\Omega}$ and estimates of the corresponding standard deviation across 1000 Monte Carlo replicates under each simulation setting specified by the design of an FG distribution and the level of $n$.
Table 1: Frequentist and Bayesian inference results in experiment (E1) across 1000 Monte Carlo replicates. Here, point.est stands for the average of 1000 point estimates for each parameter from each method, $\widehat{\text{s.d.}}$ stands for the average of the corresponding 1000 estimated standard deviations, and s.d. refers to the empirical standard deviation of the 1000 point estimates from each method. Numbers in parentheses are 100×100$\times$100 × Monte Carlo standard errors associated with the averages of 1000 estimates of the standard deviation associated with a point estimator.
@@
According to Table 1, all parameter estimates in $\mathbf{\Omega}$ are reasonably close to the true values. When the sample size is as small as 50, estimates for $\mathbf{\Omega}$ resulting from the frequentist method are still similar to those from the Bayesian inference method, although estimates for the standard deviations of these point estimators can be fairly different. We do not find such discrepancy surprising because, for the frequentist method where we use the sandwich variance estimator to infer the uncertainty of an MLE for $\mathbf{\Omega}$, the asymptotic properties associated with MLEs that support the use of a sandwich variance estimator may not take effect yet at the current sample size; and, for the Bayesian method, the quantification of standard deviation can be sensitive to the choice of priors when $n$ is small. These are confirmed by the diminishing discrepancy between the two sets of standard deviation estimates when $n=100$, 200. A closer inspection of the reported empirical mean of estimates for $\mathbf{\Omega}$ along with their empirical standard error suggests that, when $n=100$, the Bayesian method may slightly underestimate $\sigma_{2}$, the larger of the two scale parameters of FG. We believe that this is due to the inverse gamma prior imposed on the scale parameters that is sharply peaked near zero, and thus the posterior median of the larger scale parameter tends to be pulled downwards when the sample size is not large. As the sample size increases to $200$, this trend of underestimation appears to diminish. The empirical means of the standard deviation estimates from both methods are close to the corresponding empirical standard deviations, which indicate that the variability of a point estimator is accurately estimated when $n$ is not small, whether it is based on the sandwich variance estimator in the frequentist framework, or based on the posterior sampling in the Bayesian framework. In summary, the methods proposed in Section 3 under both frameworks provide reliable inference for $\mathbf{\Omega}$ along with accurate uncertainty assessment of the point estimators when data arise from an FG distribution.
@@
Among all existing mixture distributions, normal mixtures probably have the longest history and are most referenced in the literature. In another experiment, we compared the model fitting of normal mixture with that of FG when data arise from three heavy-tailed distributions: (E2) Laplace with the location parameter equal to zero and the scale parameter equal to 2; (E3) a mixture of two Gumbel distributions for the maximum, with a common mode at zero, scale parameters in the two components equal to 2 and 6, respectively, and the mixing proportion equal to 0.5; (E4) a Student-$t$ distribution with degrees of freedom equal to 5. From each of the three distributions in (E2)–(E4), we generated a random sample of size $n=200$, following which we fit a two-component normal mixture model via the EM algorithm implemented using the R package mixtools, and also fit an FG model via the two algorithms described in Section 3. This model fitting exercise was repeated for 1000 Monte Carlo replicates under each of (E2)–(E4).
@@
We used an empirical version of the Kullback-Leibler divergence as the metric to assess the quality of modeling fitting. We denote the true density function as $p(\cdot)$, and let $\hat{p}(\cdot)$ be a generic estimated density resulting from one of the three considered model fitting strategies. Under each setting in (E2)–(E4), a random sample of size 50000, $(x_{1},\ldots,x_{50000})$, was generated from the true distribution, and an empirical version of the Kullback-Leibler divergence from $\hat{p}(\cdot)$ to $p(\cdot)$ is given by $D_{\hbox{\mbox{KL}}}=(1/50000)\sum_{i=1}^{50000}\log(p(x_{i})/\hat{p}(x_{i}))$. Figure 1 shows the boxplots of $D_{\hbox{\mbox{KL}}}$ across 1000 Monte Carlo replicates corresponding to each model fitting scheme under (E2)–(E4).
@@
Judging from Figure 1, the FG distribution clearly outperforms the normal mixture when fitting data from any of the three heavy-tailed distributions in (E2)–(E4), and results from the frequentist method are comparable with those from the Bayesian method for fitting an FG model. When implementing the ECM algorithm for fitting the FG model and the EM algorithm for fitting the normal mixture, we set a maximum number of iterations at 1000. Our ECM algorithm always converged in the simulation, i.e., converged to a stationary point within 1000 iterations. However, the EM algorithm for fitting a normal mixture often had trouble achieving that, with more difficulty when data come from a heavier-tailed distribution. More specifically, under (E4), which has the highest kurtosis (equal to 9) among the three settings, the EM algorithm failed to converge in 59.9% of all Monte Carlo replicates; under (E2), which has the second highest kurtosis (equal to 6), it failed to converge in 6.7% of the replicates. Results associated with the normal mixture from these failing replicates were not included when producing the boxplots in Figure 1. In conclusion, the FG distribution is more suitable for symmetric or asymmetric heavy-tailed data than the normal mixture distribution.
@@ 5 An application in hydrology
Daily maximum water elevation changes of a water body, such as ocean, lake, and wetland, are of interest in hydrologic research. These changes may be close to zero on most days but can be extremely large or small under extreme weather. From the National Water Information System (https://waterdata.usgs.gov/), we downloaded water elevation data for Lake Murray near Columbia, South Carolina, United States, recorded from September 18, 2020 to September 18, 2021. The water elevation change of a given day was calculated by contrasting the maximum elevation and the minimum elevation on that day, returning a positive (negative) value if the maximum record of the day comes after (before) the minimum record on the same day. We fit the FG distribution to the resultant data with $n=366$ records using the frequentist method and the Bayesian method, with results presented in Table 2. The two inference methods produced very similar estimates for most parameters, although small differences were observed. For example, one would estimate the mode of daily maximum water elevation change to be $-0.795$ feet based on the frequentist method, but estimate it to be $-0.485$ feet using the Bayesian method. The discrepancy between these two mode estimates is minimal considering that the daily maximum water elevation changes range from $-38$ feet to 49.4 feet within this one year. Taking into account the uncertainty in these point estimates, we do not interpret any of these differences as statistically significant because a parameter estimate from one method always falls in the interval estimate for the same parameter from the other method according to Table 2. Using parameter estimates in Table 2 in the aforementioned R Shiny app, we obtained an estimated skewness of $-0.102$ and an estimated kurtosis of 6.384 based on the frequentist inference results, whereas the Bayes inference yielded an estimated skewness of 0.058 and an estimated kurtosis of 6.074. Combining these two sets of results, we concluded that the underlying distribution of daily maximum water elevation change may be nearly symmetrical, with outliers on both tails that cause tails heavier than that of a Gumbel distribution.
@@
$\theta$0.795 0.796 -2.355 0.764 -0.485 0.695 -1.670 0.979
$\sigma_{1}$ 5.186 0.541 4.124 6.247 5.400 0.655 4.520 6.910
$\sigma_{2}$ 6.237 1.735 2.836 9.638 5.733 1.036 4.390 8.030
$w$ 0.698 0.169 0.367 1.029 0.629 0.141 0.327 0.846
Figure 2 presents the estimated density functions from these two methods, in contrast with the estimated density curve resulting from fitting the data to a two-component normal mixture, and a kernel density estimate using a Gaussian kernel with the bandwidth selected according to the method proposed by Sheather and Jones [1991]. The last estimate is fully nonparametric and served as a benchmark against which the other three density estimates were assessed graphically. The kernel density estimate is more flexible at describing varying tail behaviors, but such flexibility comes at the cost of statistical efficiency and interpretability. With the wiggly tails evident in Figure 2 for this estimate, we suspected a certain level of overfitting of the kernel density estimate. This often happens to kernel-based estimation of a function around a region where data are scarce, with a bandwidth not large enough for the region. Between the two FG density estimates, the difference is almost negligible. They both track the kernel density estimate closely over a wide range of the support around the mode. The mode of the estimated normal mixture density is close to the other three mode estimates, but the tails are much lighter than those of the other three estimated densities.
@@
Besides comparing the three parametric density estimates pictorially, we also used the Monte-Carlo-based one-sample Kolmogorov–Smirnov test to assess the goodness of fit. The $p$-values from this test are 0.223, 0.312, and 0.106 for the frequentist FG density estimate, the Bayesian FG density estimate, and the estimated normal mixture density, respectively. Although none of the $p$-values are low enough to indicate a lack of fit (at significance level 0.05 for example), the $p$-value associated with the normal mixture is much lower than those for FG. Hence, between the two null hypotheses, with one assuming an FG distribution and the other claiming a normal mixture for this data set, we find even weaker evidence to reject the former than data evidence against the latter. It is also worth noting that the Kolmogorov–Smirnov test is known to have low power to detect deviations from a posited distribution that occur in the tails [Mason and Schuenemeyer, 1983]. This may explain the above-0.05 $p$-value for the normal mixture fit of the data even though the tail of this posited distribution may be too thin for the current data. Finally, as suggested by a referee, we computed the Akaike information criterion (AIC) and the Bayesian information criterion (BIC) after fitting the FG distribution and the normal mixture distribution to the data. When assuming an FG distribution, we obtained an AIC/BIC of 2506.028/2521.638 from the frequentist method, and 2506.299/2521.909 from the Bayesian method. When assuming a mixture normal, we found the values of AIC and BIC to be 2499.821 and 2519.334, respectively. Even though the fitted normal mixture distribution produces a lower AIC/BIC than the fitted FG distribution, we argue that these metrics focus more on the overall goodness of fit, and can be more forgiving when it comes to a relatively poor fit for certain feature of a distribution, such as the tail behavior.
We used Stan to implement the Bayesian inference for the Lake Murray data. The code can be found in https://github.com/rh8liuqy/flexible_Gumbel, where the JAGS code for fitting the FG distribution is also provided. The posterior output is given in Appendix C. The output provided there indicates that our MCMC chain has converged (see the Rhat statistics).
@@ 6 An application in criminology
$\color[rgb]{0,0,0}\sigma_{1}$ 1.600 0.180 1.247 1.954 1.690 0.214 1.206 2.686
$\color[rgb]{0,0,0}\sigma_{2}$ 51.882 45.034 -36.384 140.148 19.3 19.300 0.187 133.047
Table 4: Mean regression model based on the normal distribution fitted to the crime data. Besides parameter estimates (under point.est) and the estimated standard deviations of these parameter estimates (under $\widehat{\mbox{s.d.}}$), 95% confidence intervals of the parameters are also provided (under lower 95 and upper 95).
@@
With the location parameter $\theta$ signified in the FG distribution as the mode, it is straightforward to formulate a modal regression model that explores the relationship between the response variable and predictors. To demonstrate the formulation of a modal regression model based on the FG distribution, we analyzed a data set from Agresti et al. [2021] in the area of criminology. This data set contains the percentage of college education, poverty percentage, metropolitan rate, and murder rate for the 50 states in the United States and the District of Columbia from the year 2003. The poverty percentage is the percentage of the residents with income below the poverty level; the metropolitan rate is defined as the percentage of the population living in the metropolitan area; and the murder rate is the annual number of murders per $100,000$ people in the population.
@@
Table 3 shows the inference results from the modal regression model, and Table 4 presents the inference results from the mean regression model. At $5\%$ significance level, both frequentist and Bayesian modal regression analyses confirm that there exists a negative association between the percentage of college education and the murder rate, as well as a positive association between the metropolitan rate and the murder rate. In contrast, according to the inferred mean regression model, there is a positive association between the percentage of college education and the murder rate. Such claimed positive association is intuitively difficult to justify and contradicts many published results in criminology [Hjalmarsson and Lochner, 2012, Lochner, 2020].
The scatter plot of the data in Figure 3 can shed some light on why one reaches such a drastically different conclusion on a covariate effect when mean regression is considered in place of modal regression. As shown in Figure 3, the exists an obvious outlier, the District of Columbia ( D.C.), in panels of the first row of the scatter plot matrix for instance. D.C. not only exhibited the highest murder rate but also the highest percentage of college-educated individuals. These dual characteristics position D.C., as an outlier within the dataset. Mean regression reacts to this one extreme outlier by inflating the covariate effect associated with the percentage of college education in the inferred mean regression function. Thanks to the heavy-tailed feature of the FG distribution, modal regression based on this distribution is robust to outliers, which strives to capture data features suggested by the majority of the data and is not distracted by the extreme outlier when inferring covariate effects in this application.
Lastly, to compare their overall goodness of fit for the current data, we computed AIC and BIC following fitting each regression model. Adopting the frequentist and Bayesian methods, the modal regression analysis yields AIC/BIC equal to 239.394/252.917 and 238.710/252.233, respectively. The mean regression analysis leads to AIC and BIC equal to 303.154 and 312.813, respectively. Appendix C contains the convergence diagnosis for the Bayesian inferential method applied to this data set, from which we see no concerns about convergence.
Figure 3: Scatter plot matrix of the crime data, where D.C. stands out as an extreme outlier with the highest murder rate and the highest percentage of college education.